A car is moving at uniform acceleration. Its velocity after traveling for 2 seconds is 4m/s




and is found 12 meters from where it started. If the car’s velocity at 5 seconds is 10 m/s,




determine (a) the car’s acceleration, (b) its displacement at t = 5 seconds measured




from its starting point and (c) its velocity at the start.

1
Expert's answer
2022-02-22T10:30:11-0500

The distance traveled by a car

"d=\frac{v_i+v_f}{2}t"


The initial speed

"v_i=2d/t-v_f=2*12/2-4=8\:\rm m/s"

The acceleration

"a=\frac{v_f-v_i}{t}=\frac{4-8}{2}=-2\:\rm m/s^2"

The displacement for the 5 sec

"d=v_it+at^2/2=8*5+(-2)*5^2/2=15\:\rm m"